How would I draw this out > A rectangular room is 10 longer than it is wide, and its perimeter is 64 meters, what are the length and width ?
Thank you

Answer by fractalier(6550)

Perimeter P = 2L + 2W = 64.
Here L = W + 10...now substitute and get
2(W+10) + 2W = 64
2W + 20 + 2W = 64
4W + 20 = 64
4W = 44
W = 11 m
Then the length must be
L = 21 m